OpenStudy (anonymous):

Simplify as completely as possible: 10/x+1 / 1/2 + 3/x+1

OpenStudy (anonymous):

\[\frac{ \frac{ 10 }{ x+1 } }{ 1/2 + \frac{ 3 }{ x+1} }\]

jimthompson5910 (jim_thompson5910):

what is the inner LCD in this case

OpenStudy (anonymous):

Um.. x+1? o.o Dunno

jimthompson5910 (jim_thompson5910):

close, it's actually 2(x+1)

jimthompson5910 (jim_thompson5910):

now multiply every term by this inner LCD

jimthompson5910 (jim_thompson5910):

to clear out the inner fractions

OpenStudy (anonymous):

Okay one sec

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

\frac{ \frac{ 20(10x+10 }{ 2x(x ^{2} +1)} } }{ \frac{ 6(3x+3) }{ 2x(x ^{2} +1 } }

OpenStudy (anonymous):

Oh shoot that didn't show up .

OpenStudy (anonymous):

20(10x+10)/2x(x^2+1)/1(1/2x+1/2) + 6(3x+3)/2x(x^2+1)

jimthompson5910 (jim_thompson5910):

\[\large \frac{ \frac{ 20(10x+10 }{ 2x(x ^{2} +1)} } }{ \frac{ 6(3x+3) }{ 2x(x ^{2} +1 } }\] right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Did you forget to multiply that by the 1/2 ?

jimthompson5910 (jim_thompson5910):

ok the point of multiplying every term with 2(x+1) is to cancel out all the inner denominators

jimthompson5910 (jim_thompson5910):

so 10/(x+1) times 2(x+1) leaves you with 10*2 = 20 since the x+1 terms cancel

jimthompson5910 (jim_thompson5910):

see how I'm getting that bit?

OpenStudy (anonymous):

So under the 20(10x+10) etc.. would be / 20 + 6(etc.) ?

OpenStudy (anonymous):

You there @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

not sure what you mean exactly

OpenStudy (anonymous):

Nevermind lol, can you put the so far result in result in the equation format so I can be clear? The other one didn't show up when you posted.

jimthompson5910 (jim_thompson5910):

here is how I would do it

jimthompson5910 (jim_thompson5910):

\[\large \frac{ \frac{ 10 }{ x+1 } }{ 1/2 + \frac{ 3 }{ x+1} }\] \[\large \frac{ 2(x+1)\frac{ 10 }{ x+1 } }{ 2(x+1)\frac{1}{2} + 2(x+1)\frac{ 3 }{ x+1} }\] \[\large \frac{ 2*10 }{ x+1 + 2*3 }\] \[\large \frac{ 2*10 }{ x+1 + 6 }\] \[\large \frac{ 20 }{ x+7 }\]

OpenStudy (anonymous):

So that would be the final answer?

OpenStudy (anonymous):

How come you didn't multiply the 3/x+1 by the 2(x+1) ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

i did in a way, the x+1 in the denominator of 3/(x+1) cancels out with the x+1 in 2(x+1)

jimthompson5910 (jim_thompson5910):

so 3/(x+1) times 2(x+1) = 3*2 = 6

OpenStudy (anonymous):

Oh okay!

OpenStudy (anonymous):

Oh I see now! Thanks so much ! :)

jimthompson5910 (jim_thompson5910):

glad it's all making sense now